A random variable follows the continuous uniform distribution between

30

and

120

​a) Calculate the probability below for the distribution.

​P(60less than or equals≤xless than or equals≤90​)

​b) What are the mean and standard deviation of this​ distribution?

wo neng kan jian wen ti
Answer #1

Part a)


P ( 60 < X < 90 ) = ( 60 - 90 )/ ( 120 - 30 )
P ( 60 < X < 90 ) = 0.3333

Part b)


Mean ( X ) = ( a + b )/ 2
Mean ( X ) = ( 30 + 120 )/ 2
Mean ( X ) = 75


Variance ( X ) = ( b - a )2 / 12
Variance ( X ) = ( 120 - 30 )^2 / 12
variance ( X ) = 675

Standard Deviation =

Standard Deviation = 25.9808